# A comparison between geostatistical and machine learning models for spatio-temporal prediction of PM2.5 data

### v1

@@ 3.1.2 Nearest Neighbor Gaussian Process (NNGP)
where $\mathcal{N}_{i}$ is the neighboring set for $\mbox{\bf Y}_{i}$ and $\mbox{\bf Y}_{\mathcal{N}_{i}}$ are the observations in this set. Vecchia (1988) demonstrated that approximating the full conditional with a subset of $m$ nearest neighbors provides an excellent approximation in the case of spatial data and Datta et al. (2016b) extended it to spatiotemporal data. In this study, we construct the $m$ nearest neighbors using simple neighbor selection suggested by Datta et al. (2016b) and select the $m$ nearest spatial neighbor with temporal lag 1. The NNGP is implemented with GpGp package in R (Guinness J, 2021). GpGp relies on ordering of the Gaussian process observations where the conditional distribution only conditions on a small subset of previous observations in the ordering. This results a sparse Cholesky factor of the precision matrix. After the estimation step, to predict at any unsampled location $(s_{0},t_{0})$,
@@
3.1.3 Fixed Rank Kriging (FRK)
Fixed Rank Kriging (FRK) is a spatial/spatiotemporal interpolation technique that combines the strengths of Kriging with a reduced rank approximation to address the computational challenges of large spatial datasets (Cressie and Johannesson, 2008). The Gaussian process here is modeled similarly to (1) except that $\mbox{\bf w}(s,t)$ is approximated by a linear combination of $\mathbf{K}$ basis functions $\bf{\mathbf{\phi}(s,t)=(\phi_{1}(s,t),\phi_{2}(s,t),\dots,\phi_{\mathbf{K}}(s,t))}$ and the basis function coefficients $\mbox{\bf w}^{*}=(w_{1}^{*},w_{2}^{*},\dots,w_{\mathbf{K}}^{*})$(Wikle et al. (2019),Heaton et al. (2019)):
@@ 3.1.3 Fixed Rank Kriging (FRK)
where $\phi_{ri}(s,t)$ is the $i$th spatiotemporal basis function at the $r$th resolution with basis function coefficients $w_{\mathbf{ri}}^{*}$ and the total number of basis functions is given by $\mathbf{K}=\sum\limits_{r=1}^{\mathbf{R}}\mathbf{K_{r}}$. The coefficients $\mbox{\bf w}^{*}=(w_{\mathbf{ri}}^{*}\mathrel{\mathop{\ordinarycolon}}r=1,\dots,R,i=1,\dots,\mathbf{K_{r}})$ can be modeled as spatial, temporal or spatiotemporal varying (Wikle et al., 2019). Here we consider the case where $\mbox{\bf w}^{*}$ is spatial varying, thus $\mbox{\bf w}^{*}\sim\text{MVN}(0,\Sigma_{\mbox{\bf w}^{*}}(\mbox{\boldmath$\theta$}))$ and $\Sigma_{\mbox{\bf w}^{*}}(\mbox{\boldmath$\theta$})$ is chosen to be exponential covariance function, with $\mbox{\boldmath$\theta$}=(\sigma_{s},\rho_{s},\sigma_{t}=1,\rho_{t}=0)$ in equation (2). In this study, we construct $\{\phi_{i}(s,t),i=1,\dots,\mathbf{K}\}$ by taking the tensor product of a spatial basis function with a temporal basis function (Wikle et al., 2019). Thus, the spatiotemporal basis functions $\mathbf{\phi}(s,t)=\{\phi_{st,u}\mathrel{\mathop{\ordinarycolon}}u=1,\dots,r_{s}r_{t}\}=\{\phi_{p}(s)\psi_{q}(t)\mathrel{\mathop{\ordinarycolon}}p=1,\dots,r_{s};q=1,\dots,r_{t}\}$, where $r_{s}r_{t}$ are sets of spatial and temporal basis functions respectively. In this study, we choose bisquare basis functions for both spatial and temporal with two spatial resolutions and 1 temporal resolution (Wikle et al., 2019; Heaton et al., 2019)). We used 1600 basis functions with 80 basis functions for space and 20 basis functions for time and implemented this method using the FRK package in R (Zammit-Mangion and Cressie, 2021). With parameter estimates $\hat{\mbox{\boldmath$\beta$}}$, the FRK predictor at a new location $(s_{0},t_{0})$ is given by
@@
where Y as defined in (6), $c(s_{0},t_{0})=\text{Cov}(Y(s_{0},t_{0}),\mbox{\bf Y})$ and $\text{Var}(\mbox{\bf Y})=\Sigma_{y}(\mbox{\boldmath$\theta$})$. The uncertainty is given by
$\text{Var}(\hat{Y}(s_{0},t_{0}))=\mathbf{\phi}(s_{0},t_{0})^{T}\Sigma_{\mbox{\bf w}^{*}}(\hat{\mbox{\boldmath$\theta$}})\mathbf{\phi}(s_{0},t_{0})+\sigma_{\epsilon}^{2}-c(s_{0},t_{0})^{T}\Sigma_{y}^{-1}(\hat{\mbox{\boldmath$\theta$}})c(s_{0},t_{0}).$ (13)
3.2 Non-Geostatistical Methods Non-Geo
In this section will present methods that do not account for the spatiotemporal dependence in the data. We present the commonly used regression method and several machine learning (ML) algorithms. In recent years, ML algorithms have shown promise in modeling environmental data. Unlike traditional statistical models, machine learning algorithms don’t rely on assumptions about the underlying distribution of the data and can identify complex patterns in the data and learn from them to make accurate predictions (Breiman, 2001; Reid et al., 2015; Di et al., 2016)). Furthermore, these algorithms can handle high-dimensional data and model complex non-linear relationships between variables (Athmaja et al., 2017). For the machine learning algorithm, we will discuss random forests, support vector regressions, and neural networks with a brief explanation for each algorithm.
3.2.1 Regression (Reg)
@@
$Y(s,t)=\mbox{\boldmath$\mu$}(s,t)+\mbox{\boldmath$\epsilon$}(s,t),$ (14)
where, $\mbox{\boldmath$\mu$}(s,t)=\mbox{\bf X}(s,t)\mbox{\boldmath$\beta$}$ is the mean with $\mbox{\bf X}(s,t)=[X_{1}(s,t),X_{2}(s,t),\dots,X_{p}(s,t)]^{T}$ are predictors and $\beta$ is the $p$-dimensional vector of parameters and $\epsilon$(s,t) is the white noise measurement error with mean 0 and variance $\sigma_{\mbox{\boldmath$\epsilon$}}^{2}$. In this model, we will choose $\mbox{\bf X}(s,t)$ in terms of the nearest neighbor criteria. Although this model is simple to implement, it accounts for model error and allows us to obtain prediction error variance.
3.2.2 Random Forest (RF)
@@
First, the main idea of decision trees is to make tree that predicts a regression surface $\hat{f}(X|\hat{c})=\sum_{i=1}^{M}\hat{c}_{i}I(X\in R_{i})$ for a partition of regions $\{R_{i}\mathrel{\mathop{\ordinarycolon}}i=1,\ldots,M\}$ and $\hat{c}_{i}\in\mathbb{R}$ for $i=1,\ldots,M$. In each decision tree, a random sample of $m$ predictors is chosen as split candidates from the full set of $p$ predictors (Breiman, 2001). The algorithm uses the method of least squares to minimize $RSS(c)=\sum_{i=1}^{N}(y_{i}-f(x_{i}|c))^{2}$. It results in $\hat{c}_{m}=\frac{1}{N_{m}}\sum_{i\mathrel{\mathop{\ordinarycolon}}x_{i}\in R_{m}}y_{i}$ where $N_{m}=|\{i\mathrel{\mathop{\ordinarycolon}}x_{i}\in R_{m}\}|$. The $\{R_{i}\mathrel{\mathop{\ordinarycolon}}i=1,\ldots,M\}$ is found using a greedy algorithm to grow the regression tree top-down. Decision trees method are not always a strong ML method and can be improved using ensembles.
Random forests are non-parametric ensemble machine learning algorithms that require the selection of two key parameters: the number of predictors in the random subset of each node ($m$) as the default value and the number of decision trees in the forest ($n$). This algorithm constructs multiple decision trees using bootstrapped training samples. At each node, a random sample of $m$ features from the total $p$ features is selected, and the algorithm identifies the optimal feature for creating a split. Predictions from the $n$ trees are then aggregated using the mean to determine the final output. The error rate is assessed using predictions of out-of-bag samples (Ziegel, 2003). In this study, we used 500 trees to minimize the out-of-bag error , confirmed through five-fold cross-validation. Prediction coverage was evaluated by by calculating the 2.5% and 97.5% quantiles across all 500 trees. This was implemented the using the randomForest package in R (Liaw et al., 2002).
3.2.3 Support Vector Regression (SVR)
The groundwork for Support Vector Machines was provided by Vapnik and Chervonenkis (1974) and Vapnik (1999). A regression version of SVM, named Support Vector Regression (SVR), was introduced in Drucker et al. (1997). Based on selected kernel $k(x,y)$, a predictive model with input features $\mathbf{x}$ is given by
@@ 3.2.4 Neural Network (NN)
with $Z^{(0)}\mathrel{\mathop{\ordinarycolon}}=X$ and $f_{k}\mathrel{\mathop{\ordinarycolon}}X\mapsto Y_{k}$ is a nonlinear transformation of $X$ into $Y$. Also, $\Theta$ is a vector of all parameters of the neural network that include the weights/entries in the matrices and the biases of the linear maps and the parameters of the activation and output functions. An optimization algorithm is used to minimize a loss function $\ell(\Theta)=\frac{1}{K}\sum_{i=1}^{K}\ell_{i}(y_{k},f_{k}(X|\Theta))$. The loss function $\ell_{i}$ is usually the mean squared error $\ell_{i}(x,y)=\sum_{j=1}^{n}(x_{j}-y_{j})^{2}$ in case of regression problems and cross-entropy $\ell_{i}(x,y)=-\sum_{j=1}^{n}x_{j}\log(y_{j})$ in case of classification problems. An $L_{p}$- regularization is also used in which the objective function is $\ell(\Theta)+\lambda\lVert\Theta\rVert_{p}^{p}$ for $p=1,2$. The hyperparameter is tuned using cross-validation. The most common algorithm to find the optimal solution $\Theta$ is using back-propagation algorithm with the stochastic descent method.
There are different types of activation functions that can be used to activate the neurons
@@
  • Rectified Linear Unit (ReLU): $\sigma(x)=ReLU(x)=max(x,0)=x_{+}$
@@
There are two main possible output functions $g_{k}$
  • $k^{th}$ element “identity” function: $g_{k}(T)=T_{k}$ which is used for regression problems
  • Softmax function: $g_{k}(T)=\frac{e^{T_{k}}}{\sum_{\ell=1}^{K}e^{T_{\ell}}}$ which is used for classification problems.
Figure 4: Illustration of a Multi-layer Neural Network with NN hidden layers.
Ensemble neural networks (ENN) are a type of ensemble method that use multiple neural network models, each with their own architectures, hyperparameters, and training data, to collectively make predictions. The idea behind ENNs is that by leveraging the complementary strengths and weaknesses of multiple models, we can achieve better performance than any individual model alone. There are several ways to construct an ENN. One common approach is to train several independent neural networks, called base models or base learners, on the training data and then combine their outputs through some form of averaging, in case of regression, or voting, in classification problems (Hansen and Salamon, 1990). Another approach is to use a single architecture and hyperparameters, but vary the initial weights and biases or the order in which the data is presented during training (Ziegel, 2003). In this study, we constructed two neural networks, their predictions are combined to make a final omnibus prediction. Combining the predictions from the two neural networks can be done either by averaging or by assigning weights to each prediction according to the importance of the model, called stacking. The idea behind stacking is to use the predictions of multiple base models to train a higher-level “meta-model” that can make more accurate predictions than any of the individual models alone. We implemented ENN using Keras (Chollet et al., 2015) and Sklearn (Pedregosa et al., 2011) in Python (Van Rossum and Drake Jr, 1995).

### v2

@@ 3.2.2 Nearest Neighbor Gaussian Process
where $\mathcal{N}_{i}$ is the neighboring set for $\mbox{\bf Y}_{i}$ and $\mbox{\bf Y}_{\mathcal{N}_{i}}$ are the observations in this set. Vecchia (1988) demonstrated that approximating the full conditional with a subset of $m$ nearest neighbors provides an excellent approximation in the case of spatial data and Datta et al. (2016b) extended it to spatiotemporal data. In this study, we construct the $m$ nearest neighbors using simple neighbor selection suggested by Datta et al. (2016b) and select the $m$ nearest spatial neighbor with temporal lag 1. The NNGP is implemented with GpGp package in R (Guinness et al., 2021). GpGp relies on ordering of the Gaussian process observations where the conditional distribution only conditions on a small subset of previous observations in the ordering. This results a sparse Cholesky factor of the precision matrix. After the estimation step, to predict at any unsampled location $(s_{0},t_{0})$,
@@
3.2.3 Fixed Rank Kriging
Fixed rank kriging is a spatial/spatiotemporal interpolation technique that combines the strengths of kriging with a reduced rank approximation to address the computational challenges of large spatial datasets (Cressie and Johannesson, 2008). The Gaussian process here is modeled similarly to (1) except that $\mbox{\bf w}(s,t)$ is approximated by a linear combination of $\mathbf{K}$ basis functions $\bf{\mathbf{\phi}(s,t)=(\phi_{1}(s,t),\phi_{2}(s,t),\dots,\phi_{\mathbf{K}}(s,t))}$ and the basis function coefficients $\mbox{\bf w}^{*}=(w_{1}^{*},w_{2}^{*},\dots,w_{\mathbf{K}}^{*})$(Wikle et al., 2019; Heaton et al., 2019):
@@ 3.2.3 Fixed Rank Kriging
where $\phi_{ri}(s,t)$ is the $i$th spatiotemporal basis function at the $r$th resolution with basis function coefficients $w_{\mathbf{ri}}^{*}$ and the total number of basis functions is given by $\mathbf{K}=\sum\limits_{r=1}^{\mathbf{R}}\mathbf{K_{r}}$. The coefficients $\mbox{\bf w}^{*}=(w_{\mathbf{ri}}^{*}\mathrel{\mathop{\ordinarycolon}}r=1,\dots,R,i=1,\dots,\mathbf{K_{r}})$ can be modeled as spatial, temporal or spatiotemporal varying (Wikle et al., 2019). Here we consider the case where $\mbox{\bf w}^{*}$ is spatial varying, thus $\mbox{\bf w}^{*}\sim\text{MVN}(0,\Sigma_{\mbox{\bf w}^{*}}(\mbox{$\theta$}))$ and $\Sigma_{\mbox{\bf w}^{*}}(\mbox{$\theta$})$ is chosen to be exponential covariance function, with $\mbox{$\theta$}=(\sigma_{s},\rho_{s},\sigma_{t}=1,\rho_{t}=0)$ in equation (2). In this study, we construct $\{\phi_{i}(s,t),i=1,\dots,\mathbf{K}\}$ by taking the tensor product of a spatial basis function with a temporal basis function (Wikle et al., 2019). Thus, the spatiotemporal basis functions $\mathbf{\phi}(s,t)=\{\phi_{st,u}\mathrel{\mathop{\ordinarycolon}}u=1,\dots,r_{s}r_{t}\}=\{\phi_{p}(s)\psi_{q}(t)\mathrel{\mathop{\ordinarycolon}}p=1,\dots,r_{s};q=1,\dots,r_{t}\}$, where $r_{s}r_{t}$ are sets of spatial and temporal basis functions respectively. In this study, we choose bisquare basis functions for both spatial and temporal, we construct $\phi_{ri}(s,t)$} as tensor products of bisquare spatial and temporal bases placed with a space-filling design at $R=2$ spatial resolutions and a single temporal resolution. Spatial radii bracket the empirical correlation range; temporal knots target seasonal variability. The resulting $K=\sum_{r=1}^{R}K_{r}=1{,}600$ ($80$ spatial $\times$ 2020 temporal) yields a reduced-rank FRK with $K\ll N$. We conducted a sensitivity analysis over $K\in{800,1200,1600,2000}$, one vs. two spatial resolutions, and temporal knots 10,20,30{10,20,30} using rolling time-blocked cross-validation. Predictive accuracy, measured by the root mean squared error (RMSE), mean absolute deviation (MAD), and interval calibration varied only marginally across settings, with the two-resolutions, K=1,600K=1{,}600, 20-knot configuration offering the best accuracy–cost balance. With parameter estimates $\hat{\mbox{$\beta$}}$, the FRK predictor at a new location $(s_{0},t_{0})$ is given by
@@
where Y as defined in (6), $c(s_{0},t_{0})=\text{Cov}(Y(s_{0},t_{0}),\mbox{\bf Y})$ and $\text{Var}(\mbox{\bf Y})=\Sigma_{y}(\mbox{$\theta$})$. The uncertainty is given by
$\text{Var}(\hat{Y}(s_{0},t_{0}))=\mathbf{\phi}(s_{0},t_{0})^{T}\Sigma_{\mbox{\bf w}^{*}}(\hat{\mbox{$\theta$}})\mathbf{\phi}(s_{0},t_{0})+\sigma_{\epsilon}^{2}-c(s_{0},t_{0})^{T}\Sigma_{y}^{-1}(\hat{\mbox{$\theta$}})c(s_{0},t_{0}).$ (13)
We implemented this method using the FRK package in R (Zammit-Mangion and Cressie, 2021).
3.3 Non-geostatistical Methods
In this section we will present methods that do not account for the spatiotemporal dependence in the data. We present the commonly used regression method and several machine learning algorithms. In recent years, machine learning algorithms have shown promise in modeling environmental data. Unlike traditional statistical models, machine learning algorithms don’t rely on assumptions about the underlying distribution of the data and can identify complex patterns in the data and learn from them to make accurate predictions (Breiman, 2001; Reid et al., 2015; Di et al., 2016). Furthermore, these algorithms can handle high-dimensional data and model complex non-linear relationships between variables (Athmaja et al., 2017). For the machine learning algorithm, we will discuss regression, random forest, support vector regression, and neural network with a brief explanation for each algorithm.
3.3.1 Regression
@@
$Y(s,t)=\mbox{$\mu$}(s,t)+\mbox{$\epsilon$}(s,t),$ (14)
where, $\mbox{$\mu$}(s,t)=\mbox{\bf X}(s,t)\mbox{$\beta$}$ is the mean with $\mbox{\bf X}(s,t)=[X_{1}(s,t),X_{2}(s,t),\dots,X_{p}(s,t)]^{T}$ are predictors and $\beta$ is the $p$-dimensional vector of parameters and $\epsilon$(s,t) is the white noise measurement error with mean 0 and variance $\sigma_{\mbox{$\epsilon$}}^{2}$. In this model, we will choose $\mbox{\bf X}(s,t)$ in terms of the nearest neighbor criteria. Although this model is simple to implement, it accounts for model error and allows us to obtain prediction error variance.
3.3.2 Random Forest
@@
First, the main idea of decision trees is to make a tree that predicts a regression surface $\hat{f}(X|\hat{c})=\sum_{i=1}^{M}\hat{c}_{i}I(X\in R_{i})$ for a partition of regions $\{R_{i}\mathrel{\mathop{\ordinarycolon}}i=1,\ldots,M\}$ and $\hat{c}_{i}\in\mathbb{R}$ for $i=1,\ldots,M$. In each decision tree, a random sample of $m$ predictors is chosen as split candidates from the full set of $p$ predictors (Breiman, 2001). The algorithm uses the method of least squares to minimize $RSS(c)=\sum_{i=1}^{N}(y_{i}-f(x_{i}|c))^{2}$. It results in $\hat{c}_{m}=\frac{1}{N_{m}}\sum_{i\mathrel{\mathop{\ordinarycolon}}x_{i}\in R_{m}}y_{i}$ where $N_{m}=|\{i\mathrel{\mathop{\ordinarycolon}}x_{i}\in R_{m}\}|$. The $\{R_{i}\mathrel{\mathop{\ordinarycolon}}i=1,\ldots,M\}$ is found using a greedy algorithm to grow the regression tree top-down. Decision trees method is not always a strong machine learning method and can be improved using ensembles.
Random forest is a non-parametric ensemble machine learning algorithm that requires the selection of two key parameters: the number of predictors in the random subset of each node ($m$) as the default value and the number of decision trees in the forest ($n$). This algorithm constructs multiple decision trees using bootstrapped training samples. At each node, a random sample of $m$ features from the total $p$ features is selected, and the algorithm identifies the optimal feature for creating a split. Predictions from the $n$ trees are then aggregated using the mean to determine the final output. The error rate is assessed using predictions of out-of-bag samples (Hastie et al., 2009). In this study, we used 500 trees to minimize the out-of-bag error, confirmed through five-fold cross-validation. Prediction coverage was evaluated by calculating the 2.5% and 97.5% quantiles across all 500 trees. This was implemented using the randomForest package in R (Liaw et al., 2002).
3.3.3 Support Vector Regression
The groundwork for support vector machine was provided by Vapnik and Chervonenkis (1974) and Vapnik (1999). A regression version of that, named support vector regression, was introduced in Drucker et al. (1997). Based on selected kernel $k(x,y)$, a predictive model with input features $\mathbf{x}$ is given by
@@ 3.3.4 Neural Network
with $Z^{(0)}\mathrel{\mathop{\ordinarycolon}}=X$ and $f_{k}\mathrel{\mathop{\ordinarycolon}}X\mapsto Y_{k}$ is a nonlinear transformation of $X$ into $Y$. In addition, $\Theta$ is a vector of all parameters of the neural network that include the weights/entries in the matrices and the biases of the linear maps and the parameters of the activation and output functions. An optimization algorithm is used to minimize a loss function $\ell(\Theta)=\frac{1}{K}\sum_{i=1}^{K}\ell_{i}(y_{k},f_{k}(X|\Theta))$. The loss function $\ell_{i}$ is usually the mean squared error $\ell_{i}(x,y)=\sum_{j=1}^{n}(x_{j}-y_{j})^{2}$ in case of regression problems and cross-entropy $\ell_{i}(x,y)=-\sum_{j=1}^{n}x_{j}\log(y_{j})$ in case of classification problems. An $L_{p}$- regularization is also used in which the objective function is $\ell(\Theta)+\lambda\lVert\Theta\rVert_{p}^{p}$ for $p=1,2$. The hyperparameter is tuned using cross-validation. The most common algorithm to find the optimal solution $\Theta$ is using a back-propagation algorithm with the stochastic gradient descent method.
There are different types of activation functions that can be used to activate neurons
@@
  • Rectified Linear Unit (ReLU): $\sigma(x)=ReLU(x)=max(x,0)$
@@
There are two main output functions $g_{k}$:
  • $k$th element “identity” function: $g_{k}(T)=T_{k}$, which is used for regression problems
  • Softmax function: $g_{k}(T)=\frac{e^{T_{k}}}{\sum_{\ell=1}^{K}e^{T_{\ell}}}$, which is used for classification problems.
Figure 4: Illustration of a Multi-layer Neural Network with $N$ hidden layers.
Ensemble neural networks (ENN) are a type of ensemble method that use multiple neural network models, each with their own architectures, hyperparameters, and training data, to collectively make predictions. The idea behind ENNs is that by leveraging the complementary strengths and weaknesses of multiple models, we can achieve better performance than any individual model alone. There are several ways to construct an ENN. One common approach is to train several independent neural networks, called base models or base learners, on the training data and then combine their output through some form of averaging, in the case of regression, or voting, in classification problems (Hansen and Salamon, 1990). Another approach is to use a single architecture and hyperparameters, but vary the initial weights and biases or the order in which the data are presented during training (Hastie et al., 2009). In this study, we constructed two neural networks, and their predictions are combined to make a final omnibus prediction. Combining the predictions from the two neural networks can be done either by averaging or by assigning weights to each prediction according to the importance of the model, called stacking. The idea behind stacking is to use the predictions of multiple base models to train a higher-level “meta-model” that can make more accurate predictions than any of the individual models alone. In our implementation, the ENN uses two feed-forward neural networks (Base Learner 1 (NN1), Base Learner 2 (NN2)) trained by back-propagation with a stochastic gradient–based optimizer (Adam). We implemented the ENN using Keras (Chollet and others, 2015) and Sklearn (Pedregosa et al., 2011) in Python (Van Rossum and Drake Jr, 1995). NN1 is a wide–shallow network with hidden layers (512, 128), and NN2 is a deeper–narrower network with hidden layers (256, 128, 64). All hidden layers use ReLU activations with dropout and $\ell_{2}$ weight decay and the loss is the root mean squared error. The hyperparameters were tuned through a cross-validation grid search: learning rate $\in\{10^{-4},3{\times}10^{-4},10^{-3}\}$, batch size $\in\{256,512,1024\}$, dropout $\in[0.1,0.4]$, and penalty $\ell_{2}$ $\lambda\in\{10^{-4},10^{-3}\}$. The models were trained for up to 200 epochs with early stopping and reduction of learning-rate on the plateau; the configuration that minimizes cross-validation RMSE was selected. To combine base learners we used stacking consistent with Hansen and Salamon (1990): out-of-fold predictions from NN1 and NN2 formed two meta-features and a linear meta-learner (ridge regression with non-negativity and sum-to-one normalization of the weights) learned $(w_{1},w_{2})$, yielding the final predictor $\widehat{y}_{\mathrm{ENN}}(x)=w_{1}\,\widehat{y}_{\mathrm{NN1}}(x)+w_{2}\,\widehat{y}_{\mathrm{NN2}}(x)$ with $w_{1},w_{2}\geq 0$, $w_{1}{+}w_{2}=1$. After tuning, each base model was refit on the full training set with its selected hyperparameters, and the meta-learner was applied to their predictions; random seeds were fixed for reproducibility, and we report per-fold and aggregated RMSE for NN1, NN2, and the stacked ENN alongside the learned $(w_{1},w_{2})$.